OpenStudy (anonymous):

Solve the following system of equations. Express your answer as an ordered pair in the format (a,b). 2x + 5y = 16 -5x - 2y = 2

sam (.sam.):

(-2,4)

OpenStudy (anonymous):

Sam is right, I just checked his work.

sam (.sam.):

2x+5y=16 x=(16-5y)/2 -5(16-5y)/2 -2y=2 y=4 x=-2

OpenStudy (anonymous):

yeah (-2,4) is correct :)

OpenStudy (anonymous):

Thank you :)